# TriangleMix: A Lossless and Efficient Attention Pattern for Long Context Prefilling

### v1

@@ Abstract
Large Language Models (LLMs) rely on attention mechanisms whose time complexity grows quadratically with input sequence length, creating significant computational bottlenecks during the prefilling stage. Existing static sparse attention methods typically degrade accuracy, while dynamic sparsity methods introduce additional computational overhead due to runtime sparse index estimation. To address these limitations, we propose TriangleMix, a novel training-free static attention pattern. TriangleMix employs dense attention in shallow layers and switches to a triangle-shaped sparse pattern in deeper layers. Extensive experiments demonstrate that TriangleMix reduces attention overhead by 3.7×3.7$\times$3.7 × to 15.3×15.3$\times$15.3 × in deep layers, and decreases overall Time-to-First-Token (TTFT) by 12% to 32% for sequence lengths ranging from 32K to 128K, without sacrificing model accuracy. Moreover, TriangleMix can be seamlessly integrated with dynamic sparsity methods to achieve further speedup, e.g. accelerating MInference by 19% at 128K, highlighting its potential to enhance LLM inference efficiency. Implementation at https://aka.ms/TriangleMix
TriangleMix: A Lossless and Efficient Attention Pattern for Long Context Prefilling
Zhiyuan He1, Yike Zhang2thanks: Work during internship at Microsoft. , Chengruidong Zhang1, Huiqiang Jiang1, Yuqing Yang1, Lili Qiu1 1Microsoft Research, 2Tsinghua University zhiyuhe@microsoft.com
@@ 1 Introduction
(a) TriangleMix
(b) TriangleMix combined with dynamic attention
Figure 1: TriangleMix on Llama-3.1-8B-Instruct.
@@
Large Language Models (LLMs) are capable of processing input sequences of varying lengths Grattafiori et al. (2024); Achiam et al. (2023); Yang et al. (2024), a crucial ability that supports diverse downstream tasks, such as question answering Bai et al. (2023), long-form document understanding Zhao et al. (2024), and code generation Jiang et al. (2024b). However, due to the quadratic complexity of attention mechanisms, the time-to-first-token (TTFT) significantly increases as the input context length grows. As prior research has demonstrated, attention computation has become a critical bottleneck in the prefilling stage of LLMs Jiang et al. (2024a); Lai et al. (2025).
To address this bottleneck and accelerate the prefilling stage, researchers have proposed both static and dynamic sparse attention methods. Static sparse attention methods, such as StreamingLLM Xiao et al. (2023), reduce computational complexity from $O(N^{2})$ to $O(N)$ but suffer notable performance degradation on tasks with long contexts Li et al. (2024a). On the other hand, dynamic sparsity approaches, such as MInference Jiang et al. (2024a) and FlexPrefill Lai et al. (2025), aim to maintain near-full performance by dynamically computing sparse attention indices during inference.
Dynamic sparsity methods are effective for handling extremely long contexts (e.g., beyond 128K tokens). However, they incur additional overhead due to the runtime estimation of sparse block indices. While this overhead is negligible at very long sequence lengths, it becomes non-trivial for moderately long contexts ranging from 32K to 128K tokens. We measured the average attention kernel time per layer on Llama-3.1-8B-Instruct, and the results (Table 1) show that MInference can be 2× slower than dense attention at 32K, and offers only limited speedup at 64K. Similarly, FlexPrefill delivers a modest 1.4× speedup at 32K.
Figure 2: The average gradient $\mathrm{Grad}(\bm{M},l)$ of the Middle Q-K sections, measured on Llama-3.1-8B-Instruct, shows a significant decline in deeper layers. This suggests that the Middle Q-K components in deeper layers contribute minimally and might potentially be skipped to improve efficiency.
In this work, we aim to develop a static attention pattern that effectively accelerates the prefilling stage across various context lengths. To achieve this goal, we propose a novel gradient-based method to identify the crucial sections within attention maps. Our analysis demonstrates that the middle Q-K section (Figure 2 and 3) of the attention map can be safely omitted in deeper layers without incurring performance loss. Motivated by this insight, we introduce TriangleMix, a lossless and efficient static attention pattern for accelerating long-context LLMs. As depicted in Figure 1, TriangleMix applies standard dense attention in shallow layers, and switches to a triangle-shaped sparse attention pattern in deeper layers. TriangleMix offers four major advantages:
  • It is training-free, enabling direct integration with state-of-the-art pretrained LLMs.
  • It achieves nearly lossless accuracy, preserving the full capabilities of the underlying models.
@@
  • Its static triangle-shaped sparsity eliminates the need for dynamic block index estimation and reduces the computational complexity from $O(N^{2})$ to $O(N)$.
  • It is complementary to dynamic attention approaches. Replacing dynamic attention in deep layers with the static Triangle attention further enhances acceleration, compared with dynamic sparse attention alone.
@@
We conduct extensive experiments on three long-context LLMs, including Llama-3.1-8B-Instruct Grattafiori et al. (2024), Llama-3-8B-Instruct-262K GradientAI (2024), and Qwen2.5-7B-Instruct Yang et al. (2024), using all tasks from the RULER and LongBench benchmarks Hsieh et al. (2024); Bai et al. (2023). The results demonstrate that TriangleMix maintains the same accuracy as full attention. For Llama-3.1-8B-Instruct, TriangleMix decreases overall TTFT by 12% to 32% across context lengths ranging from 32K to 128K. Moreover, it can be seamlessly combined with dynamic sparsity methods without degrading performance. For example, integrating TriangleMix with MInference decreases TTFT by 19% at 128K context length for Llama-3.1-8B-Instruct.
@@ 2 Methodology
2.1 Sparse Attention
Figure 3: Three attention sections.
Figure 4: First row: Average attention score $\mathrm{Att}(\bm{M},l)$ for the Middle and Last Q-K sections; Second row: average gradient $\mathrm{Grad}(\bm{M},l)$ for the Middle and Last Q-K sections.
Method 32K 64K 128K
Dense 44 179 750
MInference 94 152 229
FlexPrefill ($\gamma=0.95$) 32 71 220
Table 1: Averaged attention kernel time in milliseconds for one layer on Llama-3.1-8B-Instruct.
@@ 2.1 Sparse Attention
[…]
To accelerate computing, sparse attention aims to find a sparse mask matrix $\bm{M}^{\prime}$ to compute the attention output:
@@
where $|\bm{A}-\bm{A}^{\prime}|$ is expected to be sufficiently small to preserve the original model performance.
Existing sparse attention methods fall into two categories: static sparsity and dynamic sparsity. In static sparsity, the sparse mask $\bm{M}^{\prime}$ is fixed and does not depend on the input. For example, StreamingLLM shows that most attention scores accumulate on the first few sink tokens and nearby tokens in a sliding window Xiao et al. (2023). This leads to the following streaming pattern:
where $si$ is the number of sink tokens and $sl$ is the sliding window size. This pattern, also called "A-shape" in previous works Jiang et al. (2024a), is very efficient with only $O(N)$ complexity. However, it causes significant performance drops for long-context LLMs Li et al. (2024a). In contrast, dynamic sparsity methods, such as MInference Jiang et al. (2024a) and FlexPrefill Lai et al. (2025), estimate $\bm{M}^{\prime}$ at inference time based on the input through an observation window or the pooling vector of $\bm{Q}$ and $\bm{K}$.
Dynamic sparsity generally provides better accuracy and acceleration for extremely long contexts (e.g., contexts exceeding 128K tokens). However, for moderately long contexts, the online estimation introduces additional time overhead. Specifically, as demonstrated in Table 1, MInference is slower than dense attention or offers small acceleration for contexts below 64K, while FlexPrefill achieves only modest improvements (approximately 1.4x speedup for 32K context).
2.2 Attention Sections
In this paper, we aim to design a static sparse attention mask that is both efficient and effective. Although a large fraction of attention scores falls into the streaming pattern, a significant performance gap still remains between streaming sparse attention and full dense attention. This leads us to the following question:
Beyond the streaming pattern, what other parts of the attention are crucial for maintaining the end-to-end performance of long-context LLMs?
To answer this, we divide the full attention matrix into three distinct sections as illustrated in Figure 3:
  • Streaming section: includes attention sink and the sliding window;
  • Last Q-K section: covers interactions between the last part of $\bm{Q}$ and $\bm{K}$, excluding the Streaming section;
  • Middle Q-K section: consists of the remaining interactions between the middle parts of $\bm{Q}$ and $\bm{K}$.
We define the Last Q-K mask as:
[…]
Similarly, the Middle Q-K mask is defined as:
2.3 Probing Attention Importance
Firstly, we inspect the average attention scores within the two sections defined by masks $\bm{M}^{\mathrm{middle}}$ and $\bm{M}^{\mathrm{last}}$. We observe that these average attention scores are quite similar, although both are notably lower than the average score observed in the streaming section $\bm{M}^{\mathrm{streaming}}$. To achieve a more precise measurement of the relative importance of these sections, we introduce a novel gradient-based probing technique.
@@ 2.3 Probing Attention Importance
Specifically, given an input $\bm{X}_{\mathrm{input}}$ containing $N$ tokens fed into a language model, we define a probing variable $\bm{\theta}\in\mathbb{R}^{L\times N\times N}$, where $\bm{\theta}_{\ell,i,j}=1$ for all layers $\ell$ and token indices $i,j$. Here, $L$ represents the total number of layers in the model. For simplicity of notation, we omit the attention head dimension here, but our proposed method generalizes naturally to both Multi-head Attention Vaswani et al. (2017) and Grouped Query Attention Ainslie et al. (2023).
@@
The model then outputs a logit prediction $\bm{Y_{\theta}}$ for the next token, represented as a vector $\bm{Y_{\theta}}=[y_{1},y_{2},\dots,y_{\text{vocab}}]$. Our focus is on a specific logit $y_{\mathrm{gt}}$, associated with the correct label or ground truth token. For instance, the correct answer in a multiple-choice scenario or the initial token of the target sequence in a needle-in-a-haystack task.
We are particularly interested in the partial derivative $\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}$, which measures the sensitivity of the model’s output to changes in attention scores. We quantify the importance of a specific attention section by computing the mean of the derivative values within that section. Formally, we define the gradient-based importance as:
$\mathrm{Grad}(\bm{M},l)=\frac{\sum_{\bm{M}_{i,j}=1}{\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}}}{\sum_{\bm{M}_{i,j}=1}{\bm{M}_{i,j}}}$
where $\bm{M}$ denotes the binary mask representing the targeted attention section, and $\mathrm{Grad}(\bm{M},l)$ indicates the average sensitivity of $y_{\mathrm{gt}}$ to changes within this section at layer $l$.
For comparison, we also define the average attention score of the section as:
$\mathrm{Att}(\bm{M},l)=\frac{\sum_{\bm{M}_{i,j}=1}{\bm{A}_{l,i,j}}}{\sum_{\bm{M}_{i,j}=1}{\bm{M}_{i,j}}}$
@@
Empirically, we leverage a key-value retrieval task to analyze the importance of different sections of attention. The language model is tasked with retrieve the corresponding value, starting from its first generated token. $y_{gt}$ is set to be the first token of the correct value. The full task prompt is provided in Appendix A.1.
We calculate $\mathrm{Grad}(\bm{M},l)$ and $\mathrm{Att}(\bm{M},l)$ for both the middle Q-K section ($\bm{M}^{\mathrm{middle}}$) and the last Q-K section ($\bm{M}^{\mathrm{last}}$) across 100 randomly generated samples. The hyperparameters are set as follows: the number of sink tokens ($si$) is 64, the sliding window size ($sl$) is 128, and the length of the last attention section ($last$) is also set to 128. Each input sequence consists of approximately 2,000 tokens. Finally, we visualize the results for $\mathrm{Grad}(\bm{M},l)$ and $\mathrm{Att}(\bm{M},l)$ in Figure 4. For clarity, attention scores and gradient values are both summed across all heads within each layer.

### v2

# TriangleMix: Accelerating Prefilling via Decoding-time Contribution Sparsity

@@ Abstract
Large Language Models (LLMs) incur quadratic attention complexity with input length, creating a major time bottleneck in the prefilling stage. Existing acceleration methods largely exploit attention score sparsity by estimating blocks with high attention scores and applying dynamic sparse attention. In this work, we identify another untapped form of sparsity in the prefilling stage, namely decoding-time contribution sparsity, where many attention blocks exhibit nontrivial attention scores during prefilling yet contribute negligibly to subsequent decoding, as indicated by gradient-based analysis. Building on this observation, we propose TriangleMix, a training-free static attention pattern that uses dense attention in a subset of layers and switches to Triangle attention in the others. Extensive experiments show that TriangleMix preserves nearly lossless performance relative to dense attention while substantially reducing attention overhead in Triangle layers. For 128K inputs, Triangle attention achieves a $15.3\times$ speedup in attention computation, significantly exceeding the acceleration of typical dynamic sparse methods ($1.9\times$$3.4\times$). Furthermore, TriangleMix can be seamlessly combined with dynamic sparsity approaches, delivering an additional 6%–19% reduction in TTFT over using dynamic sparsity alone.
(a) Attention sections.
(b) Gradient of Middle Q-K with respect to the first generated token
Figure 1: The average gradient $\mathrm{Grad}(\bm{M},l)$ of the Middle Q-K sections, measured on three models, shows a significant sparsity in deeper layers. This suggests that the Middle Q-K components in deeper layers contribute minimally to decoding and might potentially be skipped to improve efficiency.
(a) Measure contribution based on different tokens
(b) TriangleMix on Llama-3.1-8B-Instruct
Figure 2: Left: Attention computation in certain layers exhibits contribution locality. Right: The proposed TriangleMix pattern for Llama-3.1-8B-Instruct.
@@ 1 Introduction
Large Language Models (LLMs) are capable of processing input sequences of varying lengths (Grattafiori et al., 2024; Achiam et al., 2023; Yang et al., 2024), a crucial ability that supports diverse downstream tasks, such as question answering (Bai et al., 2023), long-form document understanding (Zhao et al., 2024), and code generation (Jiang et al., 2024b). However, due to the quadratic complexity of attention mechanisms, the attention computation time significantly increases as the input context length grows. As prior research has demonstrated, attention computation has become a critical bottleneck in the prefilling stage of LLMs (Jiang et al., 2024a; Lai et al., 2025).
To mitigate this bottleneck, several dynamic sparse attention techniques have been proposed, including MInference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and XAttention (Xu et al., 2025). These methods exploit attention score sparsity to accelerate computation: the majority of entries in attention matrix have negligible values and can be safely skipped, while blocks with significant scores need to be computed. Dynamic sparsity methods first estimate which regions are likely to contain high attention scores, and then selectively perform computation on these regions to approximate full attention (Jiang et al., 2024a; Lai et al., 2025; Xu et al., 2025).
Dynamic sparsity attention is built on a simple intuition: low-score blocks can be skipped, while blocks with non-trivial scores must be computed. This raises a deeper question: Must we compute all non-trivial blocks during prefilling to preserve decoding accuracy? A natural concern is that skipping them might distort the attention distribution and degrade performance. However, our analysis reveals that the situation is subtler. We uncover a new form of sparsity in the prefilling stage, which we call decoding-time contribution sparsity. Unlike score sparsity, it reflects that certain blocks, though holding non-trivial scores, have little impact on the actual decoding process.
To rigorously evaluate the necessity of each block, we analyze how its removal affects the generation of the first token following the prompt. Gradient analysis offers a direct measure of this sensitivity. As shown in Figure 1(a), we divide causal attention into three regions: Last Q–K, Middle Q–K, and the Streaming region. Our results indicate that the Streaming and Last Q–K regions are critical to decoding, while the Middle Q–K exhibits significant layer-wise sparsity pattern (Figure 1(b)). Notably, although the attention score magnitude of Middle Q–K regions are often comparable to Last Q–K, its gradients reveal substantially lower impact in certain layers.
This sparsity arises from the phenomenon that certain parts of attention scores’ contribution has locality of influence with respect to the decoding token position. As shown in Figure 2(a), many attention scores only affect predictions within a limited temporal window after they occur. The Middle Q-K region in some layers has considerable influence on an intermediate token $o_{i}$, but contribute very little to the tokens generated after the entire prompt $o_{N},o_{N+1},$.... This distinction allows us to safely skip Middle Q-K region’s attention computation of certain layers during the prefilling phase.
Motivated by this insight, we introduce TriangleMix, a simple yet effective static attention pattern for accelerating prefilling. As depicted in Figure 2(b), TriangleMix combines dense attention in some layers with a triangle-shaped sparse attention pattern in others. This design brings four key benefits:
  • Training-free: It can be applied directly to state-of-the-art pretrained LLMs without fine-tuning.
  • Nearly lossless accuracy: Despite its simplicity, TriangleMix preserves model performance, reaching accuracy comparable to sophisticated dynamic sparsity methods.
@@
  • Efficiency: Its static triangle pattern removes the need for block index estimation, reducing complexity from $O(N^{2})$ to $O(N)$ and enabling significantly simpler and faster attention kernel than dynamic sparsity attention.
  • Complementary to dynamic attention: Replacing dynamic sparsity with Triangle attention in selected layers delivers extra acceleration while maintaining performance.
@@
We conduct extensive experiments on three long-context LLMs, including Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Llama-3-8B-Instruct-262K (GradientAI, 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024), using all tasks from the RULER and LongBench benchmarks (Hsieh et al., 2024; Bai et al., 2023). Our results show that TriangleMix preserves the accuracy of full attention while substantially improving efficiency. Specifically, with 128K-token inputs, Triangle attention achieves a $15.3\times$ speedup in attention computation, far exceeding typical dynamic sparse methods ($1.9\times$$3.4\times$). For Llama-3.1-8B-Instruct, TriangleMix reduces overall TTFT by 12%–32% across context lengths from 32K to 128K. In addition, TriangleMix integrates seamlessly with dynamic sparsity approaches, yielding a further 6% to 19% decrease in TTFT compared to dynamic sparsity alone.
@@ 2 Methodology
2.1 Probing Attention Block Contribution
@@
where $\bm{Q},\bm{K},\bm{V}$ are matrices of shape $(N,d)$, and $\bm{M}$ is a causal mask matrix of shape $(N,N)$, with entries $\bm{M}_{i,j}\in\{0,1\}$. Here, $N$ represents the number of input tokens, and $c$ is a large positive constant to ensure attention scores masked by $\bm{M}_{i,j}=0$ become effectively zero after the softmax operation. To accelerate computing, sparse attention (Jiang et al., 2024a; Lai et al., 2025; Xu et al., 2025) aims to find a sparse mask matrix $\bm{M}^{\prime}$ to compute the attention output:
@@
The mask matrix $\bm{M^{\prime}}$ can be either static or dynamic. StreamingLLM (Xiao et al., 2023) and LM-infinite (Han et al., 2024) employ similar static streaming mask, which restricts attention to a few sink tokens and a sliding window of nearby tokens. Such pattern reduces the computation complexity of attention to $O(N)$ but significantly harms the performance of long-context LLMs(Li et al., 2024a).
In contrast, dynamic masks enable dynamic sparsity attention. During inference, block indices in $\bm{M^{\prime}}$ that likely yield non-trivial scores are identified from the input, and FlashAttention (Dao et al., 2022) is applied only to these blocks. Methods such as MInference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and XAttention (Xu et al., 2025) all adopt this strategy.
A central challenge for these dynamic sparsity attention lies in efficiently and accurately estimating the block indices with meaningful attention scores. MInference categorizes attention heads into three types and applies different estimation strategies, either by observing recent query tokens or by pooling vectors from $\bm{Q}$ and $\bm{K}$. FlexPrefill adopts a similar strategy but determines the head type dynamically during inference. XAttention, on the other hand, uses the sum of anti-diagonal scores to identify blocks with large attention. Despite these differences, the common objective remains the same: reliably locating blocks with meaningful attention scores. In this paper, we move beyond block index estimation and pose a more fundamental question: Must we compute all non-trivial blocks during prefilling to preserve decoding accuracy?
To answer this, we quantitatively probe each block’s contribution relative to the generated tokens in decoding, which we call its decoding-time contribution. Our goal is to prioritize computation for blocks with higher decoding-time contribution. Since gradients naturally capture such sensitivity, we propose a gradient-based method to perform this analysis.
@@
Given an input $\bm{X}_{\mathrm{input}}$ containing $N$ tokens fed into a language model, we define a probing variable $\bm{\theta}\in\mathbb{R}^{L\times N\times N}$, where $\bm{\theta}_{\ell,i,j}=1$ for all layers $\ell$ and token indices $i,j$. Here, $L$ represents the total number of layers in the model. For simplicity of notation, we omit the attention head dimension here, but our proposed method generalizes naturally to both Multi-head Attention (Vaswani et al., 2017) and Grouped Query Attention (Ainslie et al., 2023).
@@
The model then outputs a logit prediction $\bm{Y_{\theta}}$ for the next token after the $N$th token, represented as a vector $\bm{Y_{\theta}}=[y_{1},y_{2},\dots,y_{\text{vocab}}]$. Note that it is only the output of the last query. Our focus is on a specific logit $y_{\mathrm{gt}}$, associated with the correct label or ground truth token. For instance, the correct answer in a multiple-choice scenario or the initial token of the target sequence in a needle-in-a-haystack task. We are particularly interested in the partial derivative $\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}$, which measures the sensitivity of the model’s output to changes in attention scores. We quantify the importance of a specific attention section $\bm{M}^{\prime}$ by computing the mean of the derivative values within that section. Formally, we define the gradient-based importance as:
$\mathrm{Grad}(\bm{\hat{M}},l)=\frac{\sum_{\bm{\hat{M}}_{i,j}=1}{\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}}}{\sum_{\bm{\hat{M}}_{i,j}=1}{\bm{\hat{M}}_{i,j}}}$
@@
2.2 Decoding-time Sparsity in Middle Q-K sections
Figure 3: First row: Average attention score for the Middle and Last Q-K sections; Second row: average gradient $\mathrm{Grad}(\bm{M},l)$ for the Middle and Last Q-K sections.
Under the framework proposed in Section 2.1, we conduct an initial gradient-based analysis of three predefined attention sections. As illustrated in Figure 1(a), we partition attention into the Streaming section, the Middle Q–K section, and the Last Q–K section. We evaluate these regions using a key–value retrieval task in which the language model must output the corresponding value beginning with its first generated token; we set $y_{\mathrm{gt}}$ to the first token of the correct value. See Appendix A.1 for the exact region definitions and the full prompt.
First, we observe that both the attention scores and the gradients are high in the Streaming section, consistent with prior works (Xiao et al., 2023; Han et al., 2024). For the Middle Q–K and Last Q–K sections, we summarize our findings as follows:
Slightly higher attention scores in Middle Q–K. As shown in the top row of Figure 3, the average attention scores in the Middle Q–K and Last Q–K sections are similar, with the Middle Q–K section being slightly higher. This indicates that both regions contain non-trivial attention blocks.
Lower gradients in Middle Q–K. Despite the slightly higher attention scores, the Middle Q–K section exhibits substantially smaller gradients than the Last Q–K section (The bottom row of Figure 3). This suggests that, for predicting the first token, computing blocks in the Last Q–K region is more critical.
Layer-wise sparsity. The gradients within the Middle Q–K section display clear layer-wise sparsity. For example, in Llama-3.1-8B-Instruct and Llama-3-8B-262K (Figure 1(b)), the gradients of the last 16 layers are near zero, implying that block computations in this region could be skipped with minimal impact on first-token prediction. We refer to this phenomenon as decoding-time contribution sparsity, since it is defined directly by contributions to decoded tokens. Moreover, the observation extends beyond the first token: for later generated tokens, the current Middle Q–K region becomes a subset of the future Middle Q–K region.
2.3 Perplexity Analysis
Llama-3.1-8B-Instruct Llama-3-Instruct-262k Qwen2.5-7B-Instruct
Method 1024–1152 1920–2048 1024–1152 1920–2048 1024–1152 1920–2048
Dense 8.31 7.81 8.05 7.62 8.05 7.62
Skip Q (512–1024)-K 50% layers 8.58 (+0.27) 7.82 (+0.01) 8.33 (+0.28) 7.63 (+0.01) 8.16 (+0.11) 7.64 (+0.02)
Skip Q (512–1024)-K all layers 8.76 (+0.45) 7.90 (+0.09) 8.54 (+0.49) 7.72 (+0.10) 8.54 (+0.49) 7.72 (+0.10)
Table 1: Perplexity (PPL) comparison across different models and settings.
To further validate our claim, we study the impact of removing attention computation in the middle region. We randomly sample 1,024 Wikipedia articles and take the first 2,048 tokens from each. Three settings are compared: (1) full attention, (2) skipping attention with Q indices from 512 to 1,024 across all layers, and (3) skipping the same region in only half of the layers. In both (2) and (3), the Streaming section and the attention with other Q indices are still retained. For the third setting, we select the half of layers with the lowest $\mathrm{Grad}(\bm{\hat{M}},l)$, where $\bm{\hat{M}}$ denotes the skipped attention region and the gradient measures its contribution to predicting the 2049th token. For each model, the skipped layers remain fixed across all 1,024 samples, although different models may choose different sets of layers.
Table 1 reports perplexity changes for token ranges 1024–1152 and 1920–2048. At position 1024, the skipped computation corresponds to the Last Q–K section (since Q indices 512–1024 are omitted). At positions 1920–2048, the skipped region aligns with the Middle Q–K section. We find that skipping either all or half of the layers significantly increases perplexity in the 1024–1152 range, confirming our observations in Section 2.2 that the Last Q–K section is critical. In contrast, at 1920–2048, skipping half of the layers causes only a minor increase in perplexity (0.01–0.02), though skipping all Middle Q–K computations still raises perplexity noticeably.
We visualize this phenomenon in Figure 2(a). We refer to it as the locality of influence, which describes the finite-horizon effect of certain attention scores in transformer models. Specifically, even if some attention blocks have non-trivial scores, their influence on the output distribution disappears once the decoding position $n$ exceeds $t+\Delta t$, where $t$ is the query index of the block and $\Delta t$ is a small constant (e.g., $\Delta t=128$). This indicates that many attention scores only affect predictions within a limited temporal window after they occur. As a result, they influence nearby predictions but do not affect distant tokens, particularly those beyond the prompt. This distinction allows us to safely skip computing certain attention weights during the prefilling phase without relying on the magnitude of the attention weights themselves.